OpenStudy (anonymous):

5x - 3y - 10 = 0 x + y - 7 = 0 A: {(31/8, 25/8)} B: {(25/8, 31/8)} C: {(45/2, 31/2)} D: {(31/2, 45/2)}

OpenStudy (shaik0124):

multiply 2nd equation by 5 u will get 5x+5y=35 then subtact both equations u will get option A

OpenStudy (anonymous):

solve one of the equations for y and substitute that into the other equation:\[x+y-7=0~~~\implies~~~y=-x+7\]\[5x-3y-10=0~~\implies~~~5x-3(-x+7)-10=0\]\[\implies ~~5x+3x-21-10=0~~~\implies~~~8x-31=0\]now solve for x:\[8x-31=0~~~\implies ~~~8x=31~~~\implies~~~x={31\over8}\]to solve for y just plug in x to either equations:\[x={31\over8}~~~\implies~~~y=-\left({31\over8}\right)+7\]\[\implies~~~y={25\over8}\] \[\large And~~thus...~~~\left({31\over8},{25\over8}\right)\]

OpenStudy (anonymous):

@deynnalove16 helpful? :)

OpenStudy (anonymous):

@yummydum yes that was very helpful. Thank-You!!!! (: